26th August 2004 Hoyle meeting room 10.00 - 12:30 Present: JRL, PSB, MJI, DWE, WJS, STH, JMI, EGS & RWA (after coffee) Apologies: RGM Agenda ------ 1. Actions from last meeting 2. Comments on WFAU minutes 3. Recent VDMT and JAC telecons 4. Upgrades to ING archive 5. VEGA meeting report 6. Pipeline processing IPHAS data 7. Update on CPL + QFITS document 8. APM building works 9. AOB Minutes ------- 1. Actions from last meeting MJI has copied the correct version of the 040706 minutes to the online location. STH has supplied MJI with the location of the FIRES catalogues MJI has completed generating the catalogues from the FIRES data, but there has been no progress on the comparison. Ongoing <<<< JRL have distributed the QFITS/CPL document (see later) PSB STH has received a response from Simon Dye regarding the generated MSBs from the survey heads. Only Phil Lucas (GPS) has done this. STH has a copy of this and will investigate. <<<< ALL not many more comments have come in to WJS regarding the URD (v0.7). The release date is 31 August, so if there are any further comments, they should be handed in by the end of the week. MJI has managed to find storage for APM office equipment during the building work. This is in the link between the RGO building and the Observatory building. MJI added that RWA has already moved into JRL's office and that MTB is in the process of moving out. Obs 25 (where much of the computer equipment will go will get a portable cooling unit. PSB is now happy with his bit of the Q3 plan. JRL has acquired WFCAM acceptance test data from Paul Hirst. No data (science chips) was possible from Derek Ives since it's all on a PC which is still packed away on Hawaii. JRL should look into the WFCAM acceptance test data. <<<< MJI has emailed his URD (V0.7) comments to WJS. PSB has forwarded the DIC to JPE who has passed it on to Michelle Peron. STH was unsure about the status of the RIX responses. It was suggested that MJI asks JPE if he has collated them and passed them on to Michelle Peron. <<<< MJI raised the issue of a RIX response meeting at the VDMT meeting, but there was not much of a response. He will raise the issue again, but the date will now have to be mid-September. <<<< JRL has started updating the pipeline to more gracefully deal with lack of fringe (or suitable flat) frames. This should be complete by the end of the week. Ongoing. <<<< There was a general discussion about the need for a new default fringe frame. The current one is passed its sell-by date. JMI volunteered to look through the observers logs to see if any suitable data is available (or request for some fringe frames to be taken). <<<< JRL is in the process of building in an option of using local versions of the calibration catalogues into the pipeline. Ongoing. <<<< (Moral: don't depend on external services). JMI has rack mounted the apm2 and apm3 systems. 2. Comments on WFAU minutes From the last 3 WFAU meetings, MJI was interested in the discussion regarding access to the WFCAM science archive. WFAU were concerned about who would police the policy adopted. The chosen method would be that users need to register before they get access to the data. The earlier that this gets started the better eg. with UKIDSS consortium members. A couple of unrelated topics were discussed: STH needs to allocate a subtask for the next quarter to sort out the documents needed for the UK VDFS review. It was suggested that they should all be PDF files, but that the minimum of style requirements be set. <<<< JRL will be giving a poster rather than an oral presentation at the ADASS conference (24 October). JRL should sort out the poster material soon so that Richard Sword has enough time to work on it. <<<< 3. Recent VDMT and JAC telecons From the last VDMT meeting it was suggested that PSB should look at the Exposure Time Calculator (ETC) documentation for ISAAC data so that he can write a draft document for a VISTA ETC. <<<< For the 2nd August JAC telecon (no MJI) there was mainly a short discussion about details of the pipeline between JRL and Paul Hirst. The following telecon was on 23 August. There was much discussion regarding the slow reading speed of the LTO tapes from Hawaii. At the moment they are reading at 1-2 Mbyte/s ie. one night of WFCAM data would take more than 1 day to read in. This is probably due to a blocksize=1 problem. A mismatch between LTO1 and LTO2 tape drives has been eliminated as the source of the problem, as has the tape itself. A 26 Gbyte tape has been written at Cambridge (taking 14 minutes) and will be sent to Hawaii to help identify the problem. <<<< The news on the corrector front is that AMOS has the contract to repair the corrector. They have identified vibration as the source of the problems in carrying out the necessary measurements. They should now be able to accurately polish the surface to the correct shape. This should be complete by the beginning of November when the corrector will be placed on WFCAM, but WFCAM will be taken off UKIRT shortly after this. Only a few on-sky tests will be possible during this period. It is then intended that the science commissioning gets carried out in February/March, followed by 1 month of observing (LAS). WFCAM will then be next on in the summer ready for the beginning of the UDS season. This timescale will allow CASU to sort out any problems identified in the processing of the first month of survey data. 4. Upgrades to ING archive RWA reported on various upgrades that have been carried out to the ING archive. The old system was manual and it was found that some of the data tapes were more popular than others. There was concern that these tapes might get worn out. These 6600 tapes, covering 1982-1995, have been checked and the data placed on disc. This takes up 730 Gbytes. The data since 1995 has been on 11,300 CDs and 1800 DVDs. (Unclear if this is all on disc). The data stored on disc has been Rice compressed and the FITS headers corrected. The total size of the archive is 4.6 Tbytes. A problem has been identified with some of the CDs where over a 2-3 month period the ING CD writer was having difficulties and the CDs are faulty. Once these have been identified, replacements can be requested. JMI has altered the archive manager so that it is now much more efficient. The calibration frames are now automatically identified rather than RWA having to do this manually. It is hoped that a future upgrade will include reducing the data on the fly. (I think I have got most of the points covered by Bob. There was some mention about LTO tapes, but I missed that bit. Once you've had a go at the text, perhaps Bob could give it a once over). 5. VEGA meeting report The current plan for Gaia is that 1.5 FTEs will be based here and will work on preparing for the photometry bid which will be going in next Easter. The final phase will have on average 12 FTEs covering the period 2006-2012 with roughly half of them in Cambridge. Cambridge and Edinburgh have expressed an interest in the reduction of any VST survey data that is successfully applied for by the UK. There will be a meeting on 16 September to sort out the UK VST proposals (mainly from the UKIDSS community). The VST will have a 3-4 year window of opportunity where this is the only optical survey instrument in the Southern hemisphere. There was some more earned value discussion. 6. Pipeline processing IPHAS data JMI reported that he had been successfully running the pipeline on the IPHAS data (a Galactic plane survey in H_alpha, r' and i'). Each week of data (3 received so far?) is about 250 Gbytes and comes on a Firewire disc. This takes about 2 days to process, of which 1 day is raw computer processing. Each MEF can be processed in about 2 minutes. This is a very good test of the pipeline and the monitoring process. The LTO2 tape drive has been swapped to a Linux box (apm6) and works about two times faster. 7. Update on CPL + QFITS document Nuria McKay (ROE) has had a look over the document and in general agrees with our comments. She has suggested a number of workarounds for some of the problems with CPL and obtained a copy of the CPL guide for us. 8. APM building works On Friday (tomorrow), Paul Aslin will have a meeting with Estate Management and the architects in order to sort out timescales. It is probable that there will be a 3 week delay (which doesn't quite take us into the New Year!). 9. AOB Mike reminded us that CASU do have 0.8 of secretarial effort. Jeannette Gilbert has been allocated to us. He also said that the amount of project oversight has been lessened slightly in that for the eScience 2 grant all that is required is that Tim Naylor be on the VISTA Project Board (?). STH said that the next VDMT meeting will be on 3rd September and asked for our monthly reports to be sent in by Tuesday (31st August). <<<< There will be a meeting to discuss proposals with the Herschel Space Observatory (far infrared and submillimetre photometry and spectroscopy) on 21-23 September. WJS will be going. WJS said that the next VISTA Project Board meeting will be on 2nd September. Continuing Actions ------------------ MJI compare FIRES project reduced data catalogues with ours MJI raise the issue of a RIX response meeting for mid-September JRL finish updating pipeline to more gracefully deal with lack of fringe (or suitable flat) frames JRL finish building in option of using local version of calibration catalogues in pipeline New Actions ----------- JRL investigate the WFCAM acceptance test data (that was obtained from Paul Hirst) STH investigate GPS MSBs (and any others obtained from Simon Dye) MJI ask JPE if he has collated the RIX responses and passed them on to Michelle Peron JMI look through the observers logs to see if any suitable data is available to form a default fringe frame (or request some more data) STH allocate a subtask for the next quarter to sort out the documents needed for the UK VDFS review JRL sort out the poster material needed for the ADASS conference PSB look at the Exposure Time Calculator (ETC) documentation for ISAAC data so that a draft document for a VISTA ETC can be written MJI Send an LTO tape written at Cambridge to Hawaii to test slow reading problem ALL send in our monthly reports to STH by Tuesday (31st August)